# Integrated Optimal Control for Fast Charging and Active Thermal Management of Lithium-Ion Batteries in Extreme Ambient Temperatures

### v1

@@ Abstract
This paper presents an integrated control strategy for fast charging and active thermal management of Lithium-ion batteries in extreme ambient temperatures. A control-oriented thermal-NDC (nonlinear double-capacitor) battery model is proposed to describe the electrical and thermal dynamics, accounting for the impact from both an active thermal source and ambient temperature. A state-feedback model predictive control algorithm is then developed for integrated fast charging and active thermal management. Numerical experiments validate the algorithm under extreme temperatures, showing that the proposed algorithm can energy-efficiently adjust the battery temperature to enhance fast charging. Additionally, an output-feedback model predictive control algorithm with an extended Kalman filter is proposed for battery charging when states are partially measurable. Numerical experiments validate the effectiveness under extreme temperatures.
@@ I Introduction
The suitability of LiBs for specific applications is limited by concerns about battery charging techniques and operating temperature. For instance, electric vertical takeoff and landing (eVTOL) aircraft typically require batteries with a discharging rate three times that of electric vehicles (EVs) during takeoff and landing, as well as four times the fast charging frequency of EVs [8]. Charging and discharging LiBs in such systems can generate significant heat, which can pose a risk to the entire system, not to mention the existing frequent fires during EV charging [9]. Moreover, as electrified systems have been deployed in wildly diversified areas, managing battery charging quickly and safely, even under extreme ambient temperatures, has become an imperative research. Therefore, this paper aims to investigate an integrated strategy for LiB fast charging and thermal management under extreme temperatures.
@@ I-A Literature Review
Optimal LiB charging relies on accurate and comprehensive modeling of battery dynamics, including electrical, thermal, and aging aspects. Two primary categories of LiB models are widely recognized: 1) Electrochemical Models: These models are derived from electrochemical principles and aim to explain the electrochemical reactions and physical phenomena occurring inside a battery cell during charging and discharging. They are typically characterized by high-order partial differential equations [10, 11]. 2) Equivalent Circuit Models (ECMs): In contrast, ECMs replicate a battery’s current-voltage characteristics by utilizing electrical circuits consisting of resistors, capacitors, and voltage sources [12], offering excellent computational efficiency and making them particularly suitable for real-time battery energy management.
@@
Among the various aspects of LiB characteristics, safety and health (or aging) during fast charging are considered the most important factors. Many studies, e.g. [16, 18, 28, 29, 30], investigate how to incorporate safety and health-related constraints into optimization frameworks. These constraints ensure that the state and input at each time instance do not violate safety and health constraints. Proper constraint formulation is crucial for developing model-based charging strategies to maintain battery safety and health.
To summarize the literature review, there is a gap in modeling LiB systems, particularly in capturing the nonlinear electro-thermal dynamics where electrical dynamics are influenced by thermal dynamics, and vice versa. Existing thermal models also lack a description of how external heat sources affect the battery beyond ambient heat convection. This paper aims to address these gaps by proposing a model that incorporates nonlinear electro-thermal dynamics and accounts for external heat sources. The modeling of external heat sources could potentially enable battery charging under extreme ambient temperatures. Additionally, the paper aims to investigate a model-based charging strategy to ensure safe and fast charging, as well as active battery thermal management, even under extreme ambient temperatures.
@@ I-B Contributions, Organization, and Notations
This paper presents an integrated control strategy for fast charging and active thermal management of LiBs in extreme ambient temperatures. A control-oriented thermal-NDC battery model is proposed to describe the electrical and thermal dynamics, accounting for the impact from both an active thermal source and ambient temperature. The thermal-NDC model enables the development of a state-feedback MPC algorithm, which integrates fast charging and active thermal management for LiBs under extreme ambient temperatures.
Numerical experiments validate that the proposed algorithm under extreme temperatures can energy-efficiently adjust the battery temperature to enhance fast charging. Several insights are revealed and summarized in Section IV to explain why the decisions made by the proposed algorithm lead to energy-efficient fast charging. The observations and explanations are consistent with the literature. Additionally, an output-feedback MPC algorithm with an extended Kalman filter (EKF) is proposed for battery charging when states are partially measurable. Numerical experiments validate the effectiveness under extreme temperatures. The contributions of this paper are summarized as follows:
@@
    Section VI concludes this paper and discusses future work.
Notations. The real number set is denoted by $\mathbb{R}$. The natural number set is denoted by $\mathbb{N}$. Let $\llbracket a,b\rrbracket$ denote a set of all integers between integers $a$ and $b$, with both ends included. For $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$, $\boldsymbol{x}\leq\boldsymbol{y}$ indicates element-wise inequality. Let $\text{col}\{\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}\}$ denote a column stack of elements $\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}$, which may be scalars, vectors or matrices, i.e. $\text{col}\{\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}\}\triangleq{\begin{bmatrix}{\boldsymbol{v}_{1}}^{\top}&\cdots&{\boldsymbol{v}_{a}}^{\top}\end{bmatrix}}^{\top}$. For a matrix $\boldsymbol{A}\in\mathbb{R}^{n\times m}$, $A[i,j]$ indicates the entry in the $i$-th row and the $j$-th column of $\boldsymbol{A}$, $i\in\llbracket 1,n\rrbracket$, $j\in\llbracket 1,m\rrbracket$. Denote $\boldsymbol{I}_{n}$ as an identity matrix in $\mathbb{R}^{n\times n}$. The exponential function is denoted by $\text{exp}(\cdot)$, i.e. $\text{exp}(x)\triangleq e^{x}$, where $x\in\mathbb{R}$. $\text{diag}(a,\ b,\ \cdots,\ c)$ denote a square real matrix, where the non-diagonal elements are zeros and the diagonal elements are $a,b,\cdots,c\in\mathbb{R}$.
@@ II Battery Electro-thermal Modeling
Figure 1: The thermal-NDC (nonlinear double-capacitor) model.
@@
$\begin{bmatrix}\dot{V}_{\mathrm{b}}(t)\\ \dot{V}_{\mathrm{s}}(t)\end{bmatrix}$ $=\boldsymbol{A}_{1}(t)\begin{bmatrix}V_{\mathrm{b}}(t)\\ V_{\mathrm{s}}(t)\end{bmatrix}+\boldsymbol{B}_{1}I(t),$ (1e)
$V(t)$ $=h(V_{\mathrm{s}}(t))+R_{\mathrm{o,T}}(t)I(t),$ (1f)
where $V_{\mathrm{b}}(t)$ and $V_{\mathrm{s}}(t)$ are the voltage across $C_{\mathrm{b}}$ and $C_{\mathrm{s}}$, respectively; $I(t)$ is the input current; $I(t)<0$ for discharging and $I(t)>0$ for charging; $h:\mathbb{R}\mapsto\mathbb{R}$ denotes a mapping from $V_{\mathrm{s}}(t)$ to the open circuit voltage source, which is determined by experiments; the matrices $A_{1}(t)$ and $B_{1}$ are given by
$\boldsymbol{A}_{1}(t)=\begin{bmatrix}\dfrac{-1}{C_{\mathrm{b}}R_{\mathrm{b,T}}(t)}&\dfrac{1}{C_{\mathrm{b}}R_{\mathrm{b,T}}(t)}\\ \dfrac{1}{C_{\mathrm{s}}R_{\mathrm{b,T}}(t)}&\dfrac{-1}{C_{\mathrm{s}}R_{\mathrm{b,T}}(t)}\end{bmatrix},\ \boldsymbol{B}_{1}=\begin{bmatrix}0\\ \dfrac{1}{C_{\mathrm{s}}}\end{bmatrix}.$ (2)
@@
$\mathrm{SoC}(t)=(C_{\mathrm{b}}V_{\mathrm{b}}(t)+C_{\mathrm{s}}V_{\mathrm{s}}(t))/(C_{\mathrm{b}}+C_{\mathrm{s}})\times 100\%.$ (3)
This paper aims to formulate a minimal state-space representation for the battery dynamics, thus SoC is not a state of the system but acts as another state coordinate of the states $V_{\mathrm{b}}$ and $V_{\mathrm{b}}$. In other words, according to (3), one can determine SoC by $V_{\mathrm{b}}$ and $V_{\mathrm{b}}$, and replace SoC by $(C_{\mathrm{b}}V_{\mathrm{b}}(t)+C_{\mathrm{s}}V_{\mathrm{s}}(t))/(C_{\mathrm{b}}+C_{\mathrm{s}})$. Here, $V_{\mathrm{b}}=V_{\mathrm{s}}=0$ V when the cell is depleted ($\mathrm{SoC}=0\%$), and $V_{\mathrm{b}}=V_{\mathrm{s}}=1$ V when the cell is fully charged ($\mathrm{SoC}=100\%$). The internal resistance $R_{\mathrm{o,T}}$ is assumed to be dependent on both temperature and SoC and is given by:
@@
where $\kappa_{1},\gamma_{1},\gamma_{2},\gamma_{3}\in\mathbb{R}$ are battery-dependent parameters and determined by experiments; $T_{\mathrm{core}}$ is the core temperature of the cell, which is later described in the lumped thermal model; $T_{\mathrm{ref}}$ is the reference temperature in the Arrhenius law. The diffusion resistance $R_{\mathrm{b,T}}$ is also temperature-dependent and given by:
@@
As shown in the upper portion of Fig. 1, a two-state lumped thermal model is used to capture the thermal dynamics of a cylindrical LiB cell. The temperature along the cell’s axial direction is assumed to be uniform. The cell’s temperature distribution along the radial direction is simplified to two singular points which represent the core and surface. The battery surface temperature is affected by $\dot{Q}_{\mathrm{act}}$ from active thermal input and the diffusion between the surface and the environment. The second portion of the thermal-NDC model’s governing equations is given by:
@@
where the mapping $\boldsymbol{f}_{\mathrm{c}}:\mathbb{R}^{4}\times\mathbb{R}^{2}\mapsto\mathbb{R}^{4}$ and $\boldsymbol{g}:\mathbb{R}^{4}\times\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$ are nonlinear and given by the combination of (II) - (8).
@@ III-A Constraints
The current, terminal voltage and temperature must be within limits, i.e. $\forall t$,
@@
$V_{\mathrm{s}}-V_{\mathrm{b}}$ indicates the Li-ion concentration gradient within the electrode, which is related to the battery’s capacity degradation and cycle life. The constraint is designed to be
@@
$\zeta(t)\leq\beta_{2},\ \forall t,$ (15)
where $\zeta(t)\triangleq-\frac{C_{\mathrm{b}}+C_{\mathrm{s}}+\beta_{1}C_{\mathrm{b}}}{C_{\mathrm{b}}+C_{\mathrm{s}}}V_{\mathrm{b}}(t)+\frac{C_{\mathrm{b}}+C_{\mathrm{s}}-\beta_{1}C_{\mathrm{s}}}{C_{\mathrm{b}}+C_{\mathrm{s}}}V_{\mathrm{s}}(t)$. Finally, a constraint on the active battery temperature regulation system is considered, i.e.

### v3

# Integrated Optimal Fast Charging and Active Thermal Management of Lithium-Ion Batteries in Extreme Ambient Temperatures

@@ Abstract
This paper presents an integrated control strategy for optimal fast charging and active thermal management of Lithium-ion batteries in extreme ambient temperatures, striking a balance between charging speed and battery health. A control-oriented thermal-NDC (nonlinear double-capacitor) battery model is proposed to describe the electrical and thermal dynamics, incorporating the effects of both an active thermal source and ambient temperature. A state-feedback model predictive control algorithm is then developed for optimal fast charging and active thermal management. Numerical experiments validate the algorithm under extreme temperatures, showing that the proposed algorithm can energy-efficiently adjust the battery temperature, thereby balancing charging speed and battery health. Additionally, an output-feedback model predictive control algorithm with an extended Kalman filter is proposed for battery charging when states are partially measurable. Numerical experiments validate the effectiveness under extreme temperatures.
@@ I Introduction
The suitability of LiBs for specific applications is limited by concerns about battery charging techniques and operating temperature. For instance, electric vertical takeoff and landing (eVTOL) aircraft typically require batteries with a discharging rate three times that of electric vehicles (EVs) during takeoff and landing, as well as four times the fast charging frequency of EVs [8]. Charging and discharging LiBs in such systems can generate significant heat, which can pose a risk to the entire system, not to mention the existing frequent fires during EV charging [9]. Moreover, as electrified systems have been deployed in wildly diversified areas, managing battery charging quickly and safely, even under extreme ambient temperatures, has become an imperative research topic. Therefore, this paper aims to investigate an integrated strategy for LiB optimal fast charging and thermal management under extreme temperatures, striking a balance between charging speed and battery health.
@@ I-A Literature Review
Optimal LiB charging relies on accurate and comprehensive modeling of battery dynamics, including electrical, thermal, and aging aspects. Two primary categories of LiB models are widely recognized. 1) Electrochemical Models: These models are derived from electrochemical principles and aim to explain the electrochemical reactions and physical phenomena occurring inside a battery cell during charging and discharging. They are typically characterized by high-order partial differential equations [10, 11]. 2) Equivalent Circuit Models (ECMs): In contrast, ECMs replicate a battery’s current-voltage characteristics by utilizing electrical circuits consisting of resistors, capacitors, and voltage sources [12], offering excellent computational efficiency and making them particularly suitable for real-time battery energy management.
@@
In addition, recent literature has been exploring several emerging techniques for real-time battery charging control. For example, a two-level architecture is proposed in [30, 31, 32]. This architecture employs a low-level linear quadratic regulator to ensure stability and fast-tracking of a reference, while a high-level reference governor [33] enforces constraint satisfaction by adjusting the reference. Feng et al. [34] apply cascade-CBF (control barrier function) to modify the charging control from the CC/CV strategy, using the same model as in [17]. This approach prevents violations of voltage, SoC, and battery surface temperature constraints. Robust MPC is utilized in [35, 36] to design optimal control strategies that prevent constraint violations under system and process uncertainties. Li et al. [37] converts a constrained optimal control problem into an output tracking problem through nonlinear model inversion-based control, thereby improving real-time computation performance.
Among the various aspects of LiB characteristics, safety and health (or aging) during fast charging are considered the most important factors. Many studies, e.g. [16, 18, 28, 29, 30, 31, 32, 34, 35, 36, 37], investigate how to incorporate safety and health-related constraints into optimization frameworks. These constraints ensure that the state and input at each time instance do not violate safety and health constraints. Proper constraint formulation is crucial for developing model-based charging strategies to maintain battery safety and health.
To summarize the literature review, there is a gap in modeling LiB systems, particularly in capturing the nonlinear electro-thermal dynamics where electrical dynamics are influenced by thermal dynamics, and vice versa. Existing thermal models also lack a description of how external heat sources affect the battery beyond ambient heat convection. This paper aims to address these gaps by proposing a model that incorporates nonlinear electro-thermal dynamics and accounts for external heat sources. The modeling of external heat sources could potentially enable battery charging under extreme ambient temperatures. Additionally, the paper aims to investigate a model-based charging strategy to ensure safe, health, and fast charging, as well as active battery thermal management, even under extreme ambient temperatures.
@@ I-B Contributions, Organization, and Notations
This paper presents an integrated control strategy for optimal fast charging and active thermal management of LiBs in extreme ambient temperatures. A control-oriented thermal-NDC battery model is proposed to describe the electrical and thermal dynamics, accounting for the impact from both an active thermal source and ambient temperature. The thermal-NDC model enables the development of a state-feedback MPC algorithm, which integrates optimal fast charging and active thermal management for LiBs under extreme ambient temperatures.
Numerical experiments validate that the proposed algorithm under extreme temperatures can energy-efficiently adjust the battery temperature, thereby balancing charging speed and battery health. Several insights are revealed and summarized in Section IV to explain why the decisions made by the proposed algorithm lead to energy-efficient and health-aware optimal fast charging. The observations and explanations are consistent with the literature. Additionally, an output-feedback MPC algorithm with an extended Kalman filter (EKF) is proposed for battery charging when states are partially measurable. Numerical experiments validate the effectiveness under extreme temperatures. The contributions of this paper are summarized as follows:
@@
    Section VI concludes this paper and discusses limitations and future work.
Notations. The real number set is denoted by $\mathbb{R}$. The natural number set is denoted by $\mathbb{N}$. Let $\llbracket a,b\rrbracket$ denote a set of all integers between integers $a$ and $b$, with both ends included. For $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$, $\boldsymbol{x}\leq\boldsymbol{y}$ indicates element-wise inequality. Let $\text{col}\{\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}\}$ denote a column stack of elements $\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}$, which may be scalars, vectors or matrices, i.e., $\text{col}\{\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{a}\}\triangleq{\begin{bmatrix}{\boldsymbol{v}_{1}}^{\top}&\cdots&{\boldsymbol{v}_{a}}^{\top}\end{bmatrix}}^{\top}$. For a matrix $\boldsymbol{A}\in\mathbb{R}^{n\times m}$, $A[i,j]$ indicates the entry in the $i$-th row and the $j$-th column of $\boldsymbol{A}$, $i\in\llbracket 1,n\rrbracket$, $j\in\llbracket 1,m\rrbracket$. Denote $\boldsymbol{I}_{n}$ as an identity matrix in $\mathbb{R}^{n\times n}$. The exponential function is denoted by $\text{exp}(\cdot)$, i.e., $\text{exp}(x)\triangleq e^{x}$, where $x\in\mathbb{R}$. $\text{diag}(a,\ b,\ \cdots,\ c)$ denote a square real matrix, where the non-diagonal elements are zeros and the diagonal elements are $a,b,\cdots,c\in\mathbb{R}$.
@@ II Battery Electro-thermal Modeling
[…]
@@
$\begin{bmatrix}\dot{V}_{\mathrm{b}}(t)\\ \dot{V}_{\mathrm{s}}(t)\end{bmatrix}$ $=\boldsymbol{A}_{1}(t)\begin{bmatrix}V_{\mathrm{b}}(t)\\ V_{\mathrm{s}}(t)\end{bmatrix}+\boldsymbol{B}_{1}I(t),$ (1a)
$V(t)$ $=h(V_{\mathrm{s}}(t))+R_{\mathrm{o,T}}(t)I(t),$ (1b)
[…]
$\boldsymbol{A}_{1}(t)=\begin{bmatrix}\dfrac{-1}{C_{\mathrm{b}}R_{\mathrm{b,T}}(t)}&\dfrac{1}{C_{\mathrm{b}}R_{\mathrm{b,T}}(t)}\\ \dfrac{1}{C_{\mathrm{s}}R_{\mathrm{b,T}}(t)}&\dfrac{-1}{C_{\mathrm{s}}R_{\mathrm{b,T}}(t)}\end{bmatrix},\ \boldsymbol{B}_{1}=\begin{bmatrix}0\\ \dfrac{1}{C_{\mathrm{s}}}\end{bmatrix};$ (2)
$V(t)$ is the terminal voltage; (1b) indicates that the relation between the open-circuit voltage $U=h(V_{\mathrm{s}})$ and the terminal voltage $V$ is affected by the internal resistance $R_{\mathrm{o,T}}(t)$, which is dependent on both temperature and SoC.
@@
$\mathrm{SoC}(t)=\frac{C_{\mathrm{b}}V_{\mathrm{b}}(t)+C_{\mathrm{s}}V_{\mathrm{s}}(t)}{C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}}\times 100\%,$ (3)
where $\overline{V}_{\mathrm{b}}=1$ V and $\overline{V}_{\mathrm{s}}=1$ V are the upper bounds of $V_{\mathrm{b}}(t)$ and $V_{\mathrm{s}}(t)$, respectively. This paper aims to formulate a minimal state-space representation for the battery dynamics, thus SoC is not a state of the system but acts as another state coordinate of the states $V_{\mathrm{b}}$ and $V_{\mathrm{s}}$. In other words, according to (3), one can determine SoC by $V_{\mathrm{b}}$ and $V_{\mathrm{s}}$, and replace SoC by (3). Here, $V_{\mathrm{b}}=V_{\mathrm{s}}=0$ V when the cell is depleted ($\mathrm{SoC}=0\%$), and $V_{\mathrm{b}}=V_{\mathrm{s}}=1$ V when the cell is fully charged ($\mathrm{SoC}=100\%$).
Remark 1.
$V_{\mathrm{b}}$ and $V_{\mathrm{s}}$ must have an upper bound as a battery has a maximum capacity, and the upper bound could be arbitrary. $C_{\mathrm{b}}$ and $C_{\mathrm{s}}$ would vary accordingly so that $C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}$ is consistent with the battery’s capacity. Without loss of generality and for simplicity of notation, $\overline{V}_{\mathrm{b}}=1$ V and $\overline{V}_{\mathrm{s}}=1$ V.
The internal resistance $R_{\mathrm{o,T}}\in\mathbb{R}$ is dependent on both temperature and SoC and is given by:
@@
where $\kappa_{1},\gamma_{1},\gamma_{2},\gamma_{3}\in\mathbb{R}$ are battery-dependent parameters and determined by experiments; $T_{\mathrm{core}}\in\mathbb{R}$ is the core temperature of the cell, which is later described in the lumped thermal model; $T_{\mathrm{ref}}\in\mathbb{R}$ is the reference temperature in the Arrhenius law. The diffusion resistance $R_{\mathrm{b,T}}\in\mathbb{R}$ is also temperature-dependent and given by:
@@
As shown in the upper portion of Fig. 1, a two-state lumped thermal model, inspired by [17], is used to capture the thermal dynamics of a cylindrical LiB cell. The temperature along the cell’s axial direction is assumed to be uniform. The cell’s temperature distribution along the radial direction is simplified to two singular points which represent the core and surface. The battery surface temperature is affected by $\dot{Q}_{\mathrm{act}}$ from active thermal input and the diffusion between the surface and the environment. The second portion of the thermal-NDC model’s governing equations is given by:
@@
where the mapping $\boldsymbol{f}_{\mathrm{c}}:\mathbb{R}^{4}\times\mathbb{R}^{2}\mapsto\mathbb{R}^{4}$ and $\boldsymbol{g}:\mathbb{R}^{4}\times\mathbb{R}^{2}\mapsto\mathbb{R}^{2}$ are nonlinear and given by the combination of (1) - (8).
@@ III-A Constraints
The current, terminal voltage and temperature must be within limits, i.e., $\forall t$,
@@
The charging power $I(t)V(t)$ can also be bounded by the battery charger’s power rating. This paper omits the charging power constraint since both $I(t)$ and $V(t)$ are bounded. The surface temperature $T_{\mathrm{surf}}$ largely depends on the ambient temperature $T_{\mathrm{amb}}$, whereas the core temperature $T_{\mathrm{core}}$ has much more impacts on the internal charging and discharging processes. $T_{\mathrm{surf}}$ can also be bounded by critical temperatures at which surface thermal stress could cause significant damage to the battery [38]. Nevertheless, it is reasonable to apply the bounds of $T_{\mathrm{core}}$ to $T_{\mathrm{surf}}$.
@@
$V_{\mathrm{s}}-V_{\mathrm{b}}$ represents the Li-ion concentration gradient within the electrode [16]. This gradient indicates the undesired Lithium plating, a major factor of battery degradation in capacity and cycle life [16, 39, 40, 41, 42]. Consequently, it is crucial to limit the Li-ion concentration gradient to prevent Lithium plating [16, 41, 42]. Furthermore, this restriction should become stricter as SoC increases, as the battery becomes more vulnerable to a large Li-ion concentration gradient at higher SoC levels [28]. Therefore, this constraint can be written as
@@
$\begin{split}\zeta(t)\triangleq&-\frac{C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}+C_{\mathrm{b}}\beta_{1}}{C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}}V_{\mathrm{b}}(t)+\\ &\frac{C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}-C_{\mathrm{s}}\beta_{1}}{C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}}}V_{\mathrm{s}}(t)\leq\beta_{2},\ \forall t.\end{split}$ (15)
Remark 3.
Lithium plating refers to the deposition of metallic Lithium on the surface of anode graphite particles [42]. One of the most important contributing factors to lithium plating is lithium-ion concentration gradients, as revealed in various studies [39, 40, 41, 42]. From the appendix of reference [16], $V_{\mathrm{s}}-V_{\mathrm{b}}$ plays a role similar to the Li-ion concentration gradient. Therefore, the constraint (14) or (15) prevents lithium plating during charging. Regarding the parameters of this constraint, the values of $\beta_{1}$ and $\beta_{2}$ are typically determined by the degree to which this constraint becomes stricter as SoC increases. These values can also be determined empirically from the anode plating overpotential trajectory over time, using experimental data obtained under various temperatures and charging current conditions [30, 32, 42, 43].
Finally, a constraint on the active battery temperature regulation system is considered, i.e.